The financial statements for Castile Products, Inc., are given below: Castile Products, Inc Balance Sheet December 31 AssetsAccount balances at the beginning of the year were: accounts receivable, $150,000; and inventory, $300,000. All sales were on5. Book value per share. (Round your answer to 2 decimal places.) Book value per share
Answer #1

No. of shares = 200000/5 = 40000 shares

1) Earnings per share = Total net earnings/No. of shares = 302120/40000 = 7.55

2) Dividend payout ratio = Dividend paid/Net Income = (3.05*40000)/302120 = 40.4%

3)Dividend yield ratio = Annual dividend/share price = 3.05/64 = 4.8%

4)Price earnings ratio = Current market price of share/ EPS = 64/7.55 = 8.5

5)Book value per share = Total equity/No. of shares = 979000/40000 = 24.48